# Denoising Task Difficulty-based Curriculum for Training Diffusion Models

### v2

@@ 4.2 Exploration on Difficulties of Denoising Tasks
Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to better understand task difficulties from a distributional perspective. The training of diffusion models implicitly involves learning the distribution of the reverse process of the corresponding SDE. To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process [18]. Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks [57].
@@
This observation may stem from the inherent low-dimensional manifold of image data. As is well-known (e.g., [50]), the image data is distributed on a relatively low-dimensional manifold with a narrow support and a highly peaked multi-modal structure. On the other hand, as Gaussian noise is iteratively added, the distribution of $\bm{x}_{t}$ approaches the independent Gaussian distribution in the ambient space. Consequently, the support of the manifold broadens and the score function becomes regular over the ambient space with increasing $t$. This nature of the unconditional distribution may cause the relative entropy from $p_{t}$ to $p_{t-1}$ to decrease with $t$, indicating that it is more difficult to accurately represent the dynamics of the reverse process at small $t$.
@@ 5 Methodology
In Section 4, we observe that denoising tasks at smaller $t$ are more difficult to learn by models. From these order of difficulties in denoising tasks, we propose the incorporation strategy of an easy-to-hard training scheme, that has demonstrated its effectiveness in curriculum literature [3, 13, 27], for improving diffusion models’ training.
@@ 5.1 Design of Curriculum Learning in Diffusion Models
The next consideration involves determining the boundaries for each cluster $l_{i}$. A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$. However, this method does not account for variations in noise levels across different timesteps. Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
For EDM [24] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed. Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values. Instead, we adapt our clustering approach to suit the log-normal characteristics by defining noise level clusters $C_{i}$. Specifically, we delineate $C_{i}=[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$, where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function (quantile function) of the Gaussian distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. This method segments the noise levels into intervals by reflecting their probabilistic distribution.
@@ 5.2 Pacing Strategy of Curriculum
To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum. Training for a fixed number of iterations for each curriculum stage is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2). However, the convergence rate of each curriculum phase varies significantly, as demonstrated in Fig. 1. Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al. [13]. Our pacing function utilizes the training loss to determine transition moments and transitions to the next stage occur when the training loss converges at the current stage. Specifically, we introduce the maximum patience iteration $\tau$, and if the loss does not improve consecutively for $\tau$, the current curriculum stage is terminated, and the subsequent curriculum stage is initiated. Here, the maximum patience is a fixed hyper-parameter, and the detailed process and overall curriculum learning procedure are outlined in Algorithm 1 and 2 in Appendix D, respectively.
@@ 6.1 Experimental Setup
For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset [24], 2) Class-conditional generation with CIFAR-10 [28] and ImageNet [7] datasets, and 3) Text-to-Image generation with MS-COCO dataset [33]. In 2) and 3) setups, we applied classifier-free guidance [19].
@@
We employed three exemplary diffusion architectures for experiments: DiT [44], which integrates latent diffusion models [48] with Transformer architectures [60] parameterized as $\epsilon$-prediction, EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [49] parameterized as $F$-prediction, and SiT [38] for score- and velocity-prediction. For the text-to-image generation, we incorporated a CLIP text encoder [46] as described in DTR [43].
@@ 6.2 Comparative Results
We quantitatively validate the effectiveness of our methods with various architectures-DiT [44], EDM [22], and SiT [38]- and tasks including unconditional, class-conditional, and text-to-image generation. Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla. Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions. In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums. Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ [24], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], illustrating its effectiveness regardless of data or model used.
@@ 6.3 Analysis
In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters. To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table 2. Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters, unlike conventional curriculum learning. While both training methods in (b2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement concerning the curriculum design as shown in (b1). In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters. Besides, with findings that utilizing SNR-clustering was more effective, clustering with the actual changes in noise levels enhanced the curriculum learning adaptability.
@@
As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance. To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig. 6. We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics. Notably, our approach achieves these improvements without requiring additional parameters or training iterations, thereby significantly saving time and computational resources.
@@
Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR [43] and MinSNR [14]. Initially, we observed that each sophisticated method yields a superior performance compared to the vanilla method. Meanwhile, as shown in Table. 4, the performance is significantly enhanced when we apply the proposed curriculum learning. Consequently, the curriculum approach proves to be compatible with previous promising methods such as loss weighting (MinSNR) and architectural enhancements (DTR), demonstrating our orthogonality with recent diffusion techniques.
@@ Appendix A Extended Related Work
In this section, we review works related to analyzing diffusion model behaviors in each timestep but not covered in detail in Section 2.1. Deja et al. [6] explore denoising during the backward diffusion process and observe that transition from denoising to generation exists in the backward process. Go et al. [11] investigate the affinity between denoising tasks, showing that temporal proximal denoising tasks exhibit higher task affinity. Then, they also observe that simultaneously learning all denoising tasks by one model suffers from negative transfer. They can achieve better performance than standard diffusion training by alleviating negative transfer. Lee et al. [31] analyze frequency characteristics according to timesteps and observe that high-frequency components are lost as timesteps increase. From this observation, they propose a multi-architecture multi-experts diffusion model, which utilizes multiple denoiser models specialized in each timestep interval but utilizes a transformer-like model as the timestep increases. From observations that smaller and larger models produce similar latent noise, Pan et al. [41] propose T-Stitch, which leverages a pre-trained smaller model at the beginning of the backward process to accelerate the sampling speed. Xu et al. [64] investigate the average trace-of-covariance of training targets according to timesteps, showing that it peaks in the intermediate timesteps, causing unstable training targets. For more stable training targets, they utilize weighted conditional scores with a reference batch.
@@ Appendix B Detailed Experimental Setups for Observation
In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, and $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution. For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture. Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 and 4,000 iterations (equivalent to processing 1 million images), respectively. This training was conducted until both loss and performance converged. Configuration of training optimizers and learning rates are the same as setups in Section 6.
To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$. Following [11], we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals. At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps. For this evaluation, we sampled 10K images using a DDPM sampler over 250 steps for DiT and SiT, and an Euler solver over 40 steps for the other models.
Vanilla
Naive
Ours
Figure A: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
Vanilla
Naive
Ours
Figure B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
@@
where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 39].

### v3

@@ 4.2 Exploration on Difficulties of Denoising Tasks
Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to better understand task difficulties from a distributional perspective. The training of diffusion models implicitly involves learning the distribution of the reverse process of the corresponding SDE. To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process (Ho et al., 2020). Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks (Song et al., 2021).
@@
This observation may stem from the inherent low-dimensional manifold of image data. As is well-known (e.g., (Ruderman, 1994)), the image data is distributed on a relatively low-dimensional manifold with a narrow support and a highly peaked multi-modal structure. On the other hand, as Gaussian noise is iteratively added, the distribution of $\bm{x}_{t}$ approaches the independent Gaussian distribution in the ambient space. Consequently, the support of the manifold broadens and the score function becomes regular over the ambient space with increasing $t$. This nature of the unconditional distribution may cause the relative entropy from $p_{t}$ to $p_{t-1}$ to decrease with $t$, indicating that it is more difficult to accurately represent the dynamics of the reverse process at small $t$. More discussion is in Appendix C.
@@ 5 Methodology
In Section 4, we observe that denoising tasks at smaller $t$ are more difficult to learn by models. From these order of difficulties in denoising tasks, we propose the incorporation strategy of an easy-to-hard training scheme, that has demonstrated its effectiveness in curriculum literature (Bengio et al., 2009, Hacohen & Weinshall, 2019, Kong et al., 2021), for improving diffusion models’ training.
@@ 5.1 Design of Curriculum Learning in Diffusion Models
The next consideration involves determining the boundaries for each cluster $l_{i}$. A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$. However, this method does not account for variations in noise levels across different timesteps. Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in (Go et al., 2023a), which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
For EDM (Karras et al., 2019) which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed. Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values. Instead, we adapt our clustering approach to suit the log-normal characteristics by defining noise level clusters $C_{i}$. Specifically, we delineate $C_{i}=[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$, where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function (quantile function) of the Gaussian distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. This method segments the noise levels into intervals by reflecting their probabilistic distribution.
@@ 5.2 Pacing Strategy of Curriculum
To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum. Training for a fixed number of iterations for each curriculum stage is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2). However, the convergence rate of each curriculum phase varies significantly, as demonstrated in Fig. 1. Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al. (Hacohen & Weinshall, 2019). Our pacing function utilizes the training loss to determine transition moments and transitions to the next stage occur when the training loss converges at the current stage. Specifically, we introduce the maximum patience iteration $\tau$, and if the loss does not improve consecutively for $\tau$, the current curriculum stage is terminated, and the subsequent curriculum stage is initiated. Here, the maximum patience is a fixed hyper-parameter, and the detailed process and overall curriculum learning procedure are outlined in Algorithm 1 and 2 in Appendix D, respectively.
@@ 6.1 Experimental Setup
For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset (Karras et al., 2019), 2) Class-conditional generation with CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, and 3) Text-to-Image generation with MS-COCO dataset (Lin et al., 2014). In 2) and 3) setups, we applied classifier-free guidance (Ho & Salimans, 2022).
@@
We employed three exemplary diffusion architectures for experiments: DiT (Peebles & Xie, 2022), which integrates latent diffusion models (Rombach et al., 2022) with Transformer architectures (Vaswani et al., 2017) parameterized as $\epsilon$-prediction, EDM (Karras et al., 2022), which focuses on pixel-level diffusion utilizing UNet-based architectures (Ronneberger et al., 2015) parameterized as $F$-prediction, and SiT (Ma et al., 2024) for score- and velocity-prediction. For the text-to-image generation, we incorporated a CLIP text encoder (Radford et al., 2021) as described in DTR (Park et al., 2024b).
@@ 6.2 Comparative Results
Table 2: Evaluating the effectiveness of curriculum learning with extended training iterations on the ImageNet 256x256 dataset using the DiT-L architecture.
[…]
DiT (Vanilla) 11.18 7.84
DiT + Ours 8.18 6.24
We quantitatively validate the effectiveness of our methods with various architectures-DiT (Peebles & Xie, 2022), EDM (Karras et al., 2022), and SiT (Ma et al., 2024)- and tasks including unconditional, class-conditional, and text-to-image generation. Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla. Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions. In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums. Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ (Karras et al., 2019), ImageNet (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009), and MS-COCO (Lin et al., 2014), illustrating its effectiveness regardless of data or model used.
Showing the results of longer training might demonstrate the robustness of our method in more extended training scenarios. We trained DiT-L/2 with 2M iterations and reported the results in Table 2. Our model consistently outperformed the baseline, demonstrating its effectiveness even in prolonged training. Therefore, our method proves to be robust and effective for longer training durations.
@@ 6.3 Analysis
In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters. To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table 3. Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach (Hacohen & Weinshall, 2019), which progresses from hard to easy clusters, unlike conventional curriculum learning. While both training methods in (b2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement concerning the curriculum design as shown in (b1). In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters. Besides, with findings that utilizing SNR-clustering was more effective, clustering with the actual changes in noise levels enhanced the curriculum learning adaptability.
@@
As demonstrated in previous works (Bengio et al., 2009, Hacohen & Weinshall, 2019), the adoption of curriculum learning can lead to faster convergence in model performance. To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig. 6. We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics. Notably, our approach achieves these improvements without requiring additional parameters or training iterations, thereby significantly saving time and computational resources.
@@
Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR (Park et al., 2024b) and MinSNR (Hang et al., 2023). Initially, we observed that each sophisticated method yields a superior performance compared to the vanilla method. Meanwhile, as shown in Table. 5, the performance is significantly enhanced when we apply the proposed curriculum learning. Consequently, the curriculum approach proves to be compatible with previous promising methods such as loss weighting (MinSNR) and architectural enhancements (DTR), demonstrating our orthogonality with recent diffusion techniques.
@@ Appendix A Extended Related Work
A.1 Analyzing Diffusion Model Behaviors in Each Timestep
In this section, we review works related to analyzing diffusion model behaviors in each timestep but not covered in detail in Section 2.1. Deja et al. (Deja et al., 2022) explore denoising during the backward diffusion process and observe that transition from denoising to generation exists in the backward process. Go et al. (Go et al., 2023a) investigate the affinity between denoising tasks, showing that temporal proximal denoising tasks exhibit higher task affinity. Then, they also observe that simultaneously learning all denoising tasks by one model suffers from negative transfer. They can achieve better performance than standard diffusion training by alleviating negative transfer. Lee et al. (Lee et al., 2023) analyze frequency characteristics according to timesteps and observe that high-frequency components are lost as timesteps increase. From this observation, they propose a multi-architecture multi-experts diffusion model, which utilizes multiple denoiser models specialized in each timestep interval but utilizes a transformer-like model as the timestep increases. From observations that smaller and larger models produce similar latent noise, Pan et al. (Pan et al., 2024) propose T-Stitch, which leverages a pre-trained smaller model at the beginning of the backward process to accelerate the sampling speed. Xu et al. (Xu et al., 2023) investigate the average trace-of-covariance of training targets according to timesteps, showing that it peaks in the intermediate timesteps, causing unstable training targets. For more stable training targets, they utilize weighted conditional scores with a reference batch.
A.2 Easy-to-hard training Strategy
Progressive distillation (Salimans & Ho, 2022) focuses on reducing the number of sampling steps by training the model to progressively skip more steps, while cascaded diffusion (Ho et al., 2022b) aims to improve sample quality by progressively increasing the image resolution during training. Both methods concentrate on altering the model’s behavior or structure to tackle specific challenges, such as efficiency or resolution enhancement. In contrast, our work identifies trends in task difficulty across timestep-wise denoising tasks and leverages these findings to propose an easy-to-hard training scheme. This training strategy directly addresses the order and structure of the learning process, optimizing task sequencing to enhance performance. This distinction emphasizes that our approach is fundamentally different from these methods, as it addresses a unique aspect of diffusion model training.
@@ Appendix B Detailed Experimental Setups for Observation
In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, and $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution. For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture. Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 iterations and 4,000 kimg iterations (equivalent to processing 1 million images), respectively. This training was conducted until both loss and performance converged. As illustrated in Fig. A, we additionally plotted the iterations of each timestep interval when their loss values start to oscillate. We measured this by counting the number of times the loss value increased after the step reached 100. As shown in the results, losses of all timestep intervals are stabilized within 20K iterations, while the lower timesteps reach this point more slowly. This also suggests that the convergence speed of lower timesteps tends to exhibit a slower regime. To examine specifically at the observation of convergence, we also analyzed the convergence speed on the ImageNet dataset. As shown in Fig. B, we obtained similar results as on the FFHQ dataset. Configuration of training optimizers and learning rates are the same as setups in Section 6.
Figure A: Converged points are plotted during training for each diffusion model $\mathrm{M}_{i}$ in SiT.
Figure B: Loss convergence plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT on ImageNet dataset.
To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$. Following (Go et al., 2023a), we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals. At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps. For this evaluation, we sampled 10K images using a DDPM sampler over 250 steps for DiT and SiT, and an Euler solver over 40 steps for the other models.
@@
where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them (Glasserman, 2004, McLeish, 2011).
We sampled 5,000 images to approximate the KL divergence, which is enough for Monte-Carlo sampling and might be no changes for larger samples. Despite the large amount of samples, the exploding appearance observed in Fig. 2 when $t$ is close to zero is due to the characteristics of the data distribution. The image data distribution has narrow support (roughly speaking, it is non-zero only within a narrow range) (Ruderman & Bialek, 1993, Karras et al., 2024). As $t$ increases, information about the original data distribution gradually diminishes with the modes in the distribution of $x_{t}$ vanishing towards zero.
Given this, when $t$ is close to zero (i.e. when the distribution of $x_{t}$ is still analogous to the original data distribution), the narrow support and the tendency to move towards zero give rise to a region where $p_{t-1}$ does not overlap with $p_{t}$. Consequently, when calculating the KL divergence $D_{KL}(p_{t-1}||p_{t})=E_{x\sim p_{t-1}}[\log(\frac{p_{t-1}({x})}{p_{t}({x})})]$, $x_{t-1}$ potentially falls outside the support of $p_{t}$, which leads to $p_{t}(x_{t-1})=0$ and numerical unstability. On the other hand, as $t$ increases, the accumulated noise broadens the support of $x$’s distribution, reducing the occurrence of zero values and stabilizing the numerical estimation.